OpenStudy (anonymous):

Please help! Find the 50th derivative of y=cos2x

OpenStudy (anonymous):

what is the 50th derivative of cosine?

OpenStudy (anonymous):

cosx

OpenStudy (anonymous):

I think a little induction from following the pattern of the first 4 or 5 derivatives will point the way.

OpenStudy (anonymous):

\[f'=-\sin(x)\] \[f''=-\cos(x)\] \[f^{(3)}=-\sin(x)\] \[f^{(4)}=\cos(x)\] divide 50 by 4, get a remainder of 2, so looks like you are at \(-\cos(x)\) but of course you have the chain rule to consider each time you take the derivative you will get another factor of 2 out front

OpenStudy (anonymous):

in other words your coefficient will be \(2^{50}\)

OpenStudy (anonymous):

Thanks!